# CDQuant: Accurate Post-training Weight Quantization of Large Pre-trained Models using Greedy Coordinate Descent

### v2

@@ Appendix D Detailed Downstream Evaluation Results
w16a16 - $19.78$ $70.3$ $69.1$ $18.06$ $54.1$ $80.13$ $85.47$ $79.51$ $79.43$ $73.95$
w3a16 OWC $5.96$ $50.17$ $24.8$ $3.15$ $45.99$ $66.84$ $74.83$ $66.98$ $72.47$ $67.56$
GPTQ $13.77$ $66.12$ $56.15$ $12.99$ $52.56$ $78.58$ $85.17$ $76.75$ $78.35$ $72.3$
CD $14.49$ $65.71$ $58.83$ $13.44$ $52.47$ $79.92$ $83.03$ $75.81$ $77.09$ $71.9$
BCD(k=2) $14.63$ $65.96$ $59.46$ $14.07$ $52.73$ $79.84$ $83.09$ $75.93$ $77.42$ $71.74$
w3a16g128 OWC $10.89$ $67.98$ $45.03$ $8.81$ $48.29$ $74.83$ $79.14$ $74.64$ $75.52$ $72.45$
GPTQ $16.12$ $69.8$ $60.82$ $14.47$ $53.16$ $79.34$ $84.07$ $77.12$ $77.86$ $73.16$
CD $15.46$ $71.62$ $62.07$ $16.19$ $52.3$ $78.79$ $84.34$ $77.02$ $78.35$ $74.82$
BCD(k=2) $16.04$ $71.04$ $62.02$ $15.65$ $51.71$ $78.32$ $84.83$ $77.04$ $77.69$ $74.03$
OWC-CD $15.62$ $71.78$ $61.19$ $14.76$ $52.99$ $79.34$ $84.1$ $76.92$ $77.42$ $75.22$
OWC-CD + CD $16.2$ $70.77$ $61.8$ $14.71$ $53.41$ $79.42$ $84.5$ $77.22$ $78.29$ $73.64$
OWC-CD + BCD(k=2) $16.7$ $71.3$ $61.87$ $14.91$ $53.84$ $79.8$ $84.37$ $77.25$ $78.45$ $74.19$
w4a16 OWC $16.81$ $71.7$ $61.75$ $16.54$ $53.33$ $78.24$ $85.17$ $77.32$ $79.33$ $71.9$
GPTQ $19$ $70.66$ $66.01$ $16.34$ $53.67$ $79.46$ $85.57$ $78.5$ $79.05$ $72.93$
CD $18.7$ $70.97$ $66.11$ $17.08$ $54.27$ $80.89$ $85.87$ $78.55$ $79.05$ $72.38$
BCD(k=2) $19.14$ $70.9$ $65.72$ $16.88$ $54.78$ $80.6$ $85.6$ $78.64$ $79.65$ $73.72$
w4a16g128 OWC $18.03$ $69.97$ $65.28$ $15.94$ $53.41$ $79.42$ $84.89$ $78.1$ $79.16$ $73.56$
GPTQ $19.17$ $70.34$ $67.46$ $17.18$ $53.58$ $79.88$ $85.02$ $79$ $79.11$ $73.56$
CD $19.25$ $71.12$ $67.61$ $17.13$ $53.33$ $80.18$ $84.92$ $79.09$ $79.11$ $72.93$
BCD(k=2) $19.45$ $70.81$ $67.75$ $17.08$ $54.27$ $80.3$ $84.74$ $79.05$ $78.94$ $72.93$
OWC-CD $19.2$ $70.58$ $67.66$ $17.22$ $53.24$ $80.51$ $85.05$ $78.84$ $79.49$ $74.74$
OWC-CD + CD $19.64$ $70.8$ $68.02$ $17.52$ $53.75$ $80.77$ $84.98$ $79.09$ $78.78$ $74.27$
OWC-CD + BCD(k=2) $19.31$ $70.41$ $67.91$ $17.37$ $53.67$ $80.3$ $84.95$ $79.07$ $78.78$ $74.43$
Table 9: Table presents downstream evaluation (one-shot) numbers of PaLM2-Gecko - for GPTQ, CD, BCD for INT3, INT4 quantization of FFN weights.
Config Method PaLM2-Gecko
w16a16 - $7.67$ $49.04$ $37.72$ $8.96$ $37.12$ $68.1$ $66.15$ $61.55$ $73.99$ $61.72$
w3a16 OWC $3.71$ $48.17$ $18.84$ $5.07$ $30.03$ $60.1$ $59.69$ $53.45$ $70.24$ $60.93$
GPTQ $4.38$ $46.91$ $20.43$ $5.95$ $31.23$ $62.84$ $63.55$ $56.51$ $71.65$ $60.46$
CD $4.63$ $48.1$ $22.79$ $7.23$ $32.59$ $63.51$ $62.14$ $57.6$ $72.09$ $60.38$
BCD(k=2) $4.52$ $48.84$ $22.09$ $6.64$ $32.51$ $64.14$ $63.24$ $57.53$ $72.2$ $61.01$
w3a16g128 OWC $4.43$ $57.48$ $20.99$ $8.42$ $32.59$ $63.09$ $62.05$ $53.45$ $71.38$ $59.59$
GPTQ $4.85$ $53.14$ $23.25$ $7.73$ $33.11$ $64.73$ $64.43$ $56.19$ $72.42$ $61.09$
CD $4.49$ $54.43$ $24.12$ $8.12$ $33.28$ $65.15$ $64.22$ $56.56$ $71.98$ $61.64$
BCD(k=2) $4.29$ $53.58$ $24.41$ $8.27$ $33.7$ $65.32$ $64.25$ $56.25$ $71.71$ $61.88$
OWC-CD $4.6$ $53.86$ $23.23$ $7.14$ $32.85$ $64.81$ $64.65$ $56.31$ $71.82$ $61.72$
OWC-CD + CD $4.52$ $52.4$ $25.07$ $7.78$ $32.76$ $64.44$ $65.11$ $56.85$ $71.93$ $61.96$
OWC-CD + BCD(k=2) $4.65$ $51.15$ $25.13$ $7.68$ $33.19$ $64.9$ $64.68$ $56.88$ $71.82$ $61.17$
w4a16 OWC $6.09$ $48.4$ $31.06$ $7.53$ $35.24$ $65.07$ $64.59$ $59.59$ $72.96$ $63.85$
GPTQ $6.81$ $50.67$ $32.5$ $8.07$ $36.43$ $65.53$ $65.05$ $60.16$ $73.29$ $63.85$
CD $6.81$ $51.07$ $31.87$ $8.66$ $35.58$ $66.92$ $64.65$ $60.22$ $73.23$ $63.14$
BCD(k=2) $6.65$ $50.86$ $31.67$ $8.91$ $35.15$ $67.21$ $64.4$ $60.31$ $73.45$ $63.22$
w4a16g128 OWC $7.12$ $47.85$ $33.58$ $9.99$ $35.58$ $67.09$ $68.38$ $60.15$ $73.5$ $62.9$
GPTQ $7.2$ $51.6$ $33.94$ $8.86$ $36.18$ $67.21$ $66.51$ $59.99$ $73.5$ $63.38$
CD $7.59$ $48.87$ $35.22$ $9.06$ $35.92$ $68.01$ $67.09$ $60.49$ $73.45$ $63.54$
BCD(k=2) $7.29$ $48.72$ $34.67$ $9.1$ $36.69$ $68.01$ $67.28$ $60.7$ $73.34$ $62.9$
OWC-CD $7.31$ $48.52$ $33.2$ $9.3$ $35.75$ $67.59$ $66.97$ $60.8$ $74.21$ $63.06$
OWC-CD + CD $7.15$ $49.15$ $34.77$ $9.4$ $36.52$ $67.8$ $66.33$ $60.74$ $73.61$ $63.61$
OWC-CD + BCD(k=2) $7.29$ $48.88$ $34.68$ $9.15$ $36.43$ $68.01$ $65.75$ $60.77$ $74.05$ $64.01$
Table 10: Table presents downstream evaluation (one-shot) numbers of PaLM2-Otter - for GPTQ, CD, BCD for INT2 quantization of FFN weights.
Method Epochs PaLM2-Otter
w2a16g128 NatualQ. SQuAD TriviaQA WebQ ARC-c ARC-e BoolQ HellaSwag PIQA WinoGrande
GPTQ - $6.93$ $64.89$ $28.36$ $7.73$ $44.28$ $71.04$ $78.81$ $68.29$ $72.69$ $67.72$
CD $1$ $8.23$ $68.12$ $32.98$ $10.14$ $43.69$ $72.43$ $76.91$ $68.11$ $72.96$ $66.46$
BCD(k=2) $1$ $8.09$ $69.2$ $32.85$ $9.99$ $43.43$ $73.02$ $77$ $68.03$ $73.07$ $66.38$
$2$ $8.2$ $69.43$ $32.98$ $9.74$ $43.52$ $73.7$ $77.03$ $68.56$ $74.1$ $67.8$
$3$ $8.31$ $69.79$ $33.03$ $9.65$ $43.52$ $73.11$ $77.09$ $68.76$ $73.07$ $67.01$
BCD(k=4) $1$ $7.87$ $69.52$ $33.24$ $9.79$ $43.17$ $72.6$ $77.83$ $68.61$ $73.88$ $66.46$
$2$ $8.31$ $69.06$ $33.03$ $9.6$ $43$ $72.6$ $76.64$ $68.51$ $73.18$ $66.3$
$3$ $8.31$ $69.8$ $32.69$ $9.5$ $42.15$ $72.43$ $77.06$ $68.34$ $73.72$ $67.88$
Table 11: Table presents downstream evaluation (one-shot) numbers of PaLM2-Bison - for GPTQ, CD, BCD for INT2 quantization of FFN weights.
Method Epochs PaLM2-Bison
w2a16g128 NatualQ. SQuAD TriviaQA WebQ ARC-c ARC-e BoolQ HellaSwag PIQA WinoGrande
GPTQ - $15.26$ $69.34$ $55.6$ $17.08$ $51.02$ $80.13$ $86.02$ $74.44$ $78.78$ $74.74$
CD $1$ $15.6$ $67.43$ $55.06$ $16.83$ $50.43$ $79.97$ $83.76$ $74.35$ $78.35$ $74.51$
BCD(k=2) $1$ $15.62$ $67.32$ $55.82$ $16.63$ $51.45$ $79.97$ $83.61$ $74.46$ $79$ $75.45$
$2$ $15.62$ $68.45$ $56.36$ $16.98$ $51.02$ $80.09$ $84.56$ $74.99$ $78.89$ $75.06$
$3$ $16.01$ $67.99$ $56.31$ $16.88$ $51.19$ $79.88$ $84.65$ $74.95$ $79.33$ $75.45$
BCD(k=4) $1$ $15.9$ $67.46$ $56.12$ $16.44$ $51.28$ $80.35$ $84.01$ $74.89$ $78.84$ $75.85$
$2$ $16.04$ $67.61$ $56.34$ $16.39$ $51.54$ $80.09$ $83.91$ $74.83$ $78.73$ $75.37$
$3$ $16.12$ $67.54$ $56.3$ $16.24$ $51.02$ $80.22$ $83.98$ $74.93$ $78.51$ $75.93$
Table 12: Table presents downstream evaluation (zero-shot) numbers for GPTQ, CD, BCD for INT3, INT4 quantization of FFN weights. Gen, Rank columns correspond to generation and ranking tasks.
Config Method PaLM2-Gecko PaLM2-Otter PaLM2-Bison
Gen. Rank Avg. Gen. Rank Avg. Gen. Rank Avg.
w16a16 - $25.85$ $61.44$ $47.2$ $44.31$ $75.43$ $62.98$ $51.08$ $78.28$ $68.00$
w3a16 OWC $15.45$ $55.73$ $39.62$ $15.06$ $59.29$ $41.60$ $39.58$ $75.45$ $61.10$
GPTQ $13.08$ $56.92$ $39.39$ $31.80$ $71.94$ $55.88$ $41.42$ $76.70$ $62.58$
CD $15.82$ $58.19$ $41.24$ $32.34$ $72.03$ $56.16$ $41.66$ $76.71$ $62.69$
BCD(k=2) $16.16$ $57.85$ $41.17$ $32.04$ $72.22$ $56.15$ $41.37$ $76.78$ $62.62$
w3a16g128 OWC $18.54$ $56.95$ $41.59$ $26.62$ $68.09$ $51.50$ $41.54$ $76.51$ $62.52$
GPTQ $17.96$ $58.30$ $42.16$ $33.43$ $72.98$ $57.16$ $43.47$ $76.65$ $63.38$
CD $16.81$ $58.10$ $41.58$ $32.73$ $72.62$ $56.66$ $42.95$ $77.03$ $63.4$
BCD(k=2) $16.33$ $57.81$ $41.22$ $32.13$ $72.67$ $56.46$ $42.67$ $76.85$ $63.18$
OWC-CD $16.73$ $57.74$ $41.34$ $32.90$ $72.40$ $56.60$ $42.23$ $77.02$ $63.11$
OWC-CD + CD $16.37$ $58.22$ $41.48$ $33.53$ $72.64$ $56.99$ $43.09$ $76.91$ $63.38$
w4a16 OWC $16.07$ $59.42$ $42.08$ $33.76$ $72.22$ $56.83$ $43.59$ $77.13$ $63.71$
GPTQ $18.77$ $59.5$ $43.21$ $34.03$ $73.53$ $57.73$ $44.16$ $77.54$ $64.19$
CD $18.20$ $59.95$ $43.25$ $35.01$ $73.42$ $58.06$ $43.97$ $77.55$ $64.12$
BCD(k=2) $17.89$ $59.64$ $42.94$ $35.10$ $73.31$ $58.02$ $43.95$ $77.77$ $64.24$
w4a16g128 OWC $19.00$ $60.42$ $43.85$ $34.52$ $72.83$ $57.5$ $44.45$ $77.71$ $64.41$
GPTQ $19.07$ $60.40$ $43.87$ $35.34$ $73.60$ $58.29$ $44.25$ $77.65$ $64.29$
CD $20.35$ $60.94$ $44.70$ $36.20$ $73.67$ $58.68$ $44.21$ $77.72$ $64.31$
BCD(k=2) $20.44$ $60.69$ $44.59$ $36.11$ $73.54$ $58.57$ $44.17$ $77.56$ $64.2$
OWC-CD $19.80$ $60.39$ $44.15$ $36.33$ $73.62$ $58.71$ $44.75$ $77.70$ $64.52$
OWC-CD + CD $20.47$ $60.85$ $44.70$ $35.98$ $73.59$ $58.54$ $44.77$ $77.52$ $64.42$
@@
Config Method PaLM2-Otter
w16a16 - $aa$ $aa$ $aa$ $aa$ $aa$ $aa$ $aa$ $aa$ $aa$ $aa$
w3a16 OWC $2.27$ $49.45$ $8.02$ $0.49$ $41.98$ $60.14$ $48.78$ $66.91$ $71.87$ $66.06$
GPTQ $9.09$ $64.53$ $46.38$ $7.19$ $49.83$ $73.99$ $81.96$ $76.87$ $78.89$ $70.09$
CD $10.72$ $64.01$ $48.03$ $6.59$ $50.77$ $76.6$ $81.01$ $75.77$ $77.97$ $70.09$
BCD(k=2) $10.55$ $64.31$ $47.25$ $6.05$ $50.85$ $76.43$ $81.41$ $75.91$ $78.45$ $70.24$
GPTQ $10.17$ $66.53$ $49.29$ $7.73$ $51.37$ $75.93$ $82.35$ $77$ $79.11$ $72.14$
CD $7.17$ $68.73$ $49.42$ $5.61$ $50.26$ $76.18$ $82.39$ $76.79$ $78.13$ $71.98$
BCD(k=2) $6.68$ $68.18$ $48.39$ $5.27$ $50.26$ $75.93$ $81.41$ $76.84$ $79$ $72.61$
OWC-CD $8.98$ $68.37$ $48.33$ $5.91$ $50.51$ $75.46$ $81.35$ $76.83$ $78.18$ $72.06$
OWC-CD + CD $9.72$ $67.87$ $49.87$ $6.64$ $51.37$ $76.68$ $81.59$ $76.78$ $77.91$ $71.51$
w4a16 OWC $11.19$ $66.67$ $49.56$ $7.63$ $50$ $73.91$ $80$ $77.36$ $79.11$ $72.93$
GPTQ $12.27$ $66.65$ $48.72$ $8.46$ $51.71$ $75.84$ $82.02$ $78.2$ $80.09$ $73.32$
CD $12.52$ $66.85$ $53.16$ $7.53$ $52.3$ $76.43$ $81.28$ $78.3$ $79.6$ $72.61$
BCD(k=2) $12.33$ $67$ $53.45$ $7.63$ $51.96$ $75.97$ $80.92$ $78.47$ $80.47$ $72.06$
GPTQ $11.5$ $65.81$ $56.36$ $7.68$ $51.71$ $76.52$ $82.08$ $78.86$ $79.27$ $73.16$
CD $11.72$ $66.85$ $57.63$ $8.61$ $51.19$ $76.6$ $82.51$ $78.89$ $79.49$ $73.32$
BCD(k=2) $11.66$ $66.73$ $57.5$ $8.56$ $51.19$ $76.3$ $82.29$ $78.88$ $79.27$ $73.32$
OWC-CD $12.27$ $66.28$ $58.21$ $8.56$ $51.62$ $76.85$ $82.32$ $78.85$ $79.49$ $72.61$
OWC-CD + CD $11.63$ $66.28$ $58.06$ $7.92$ $51.62$ $76.94$ $82.45$ $78.72$ $79.11$ $72.69$
Config Method PaLM2-Gecko
w16a16 - $aa$ $aa$ $aa$ $aa$ $aa$ $aa$ $aa$ $aa$ $aa$ $aa$
w3a16 OWC $2.55$ $47.57$ $9.81$ $1.87$ $29.69$ $56.73$ $64.8$ $53.28$ $69.59$ $60.3$
GPTQ $3.35$ $44.13$ $3.97$ $0.89$ $30.2$ $62.12$ $63.79$ $56.33$ $71.22$ $57.85$
CD $3.55$ $45.71$ $11.32$ $2.71$ $30.63$ $60.56$ $66.45$ $57.2$ $71.87$ $62.43$
BCD(k=2) $3.63$ $47.04$ $11.22$ $2.76$ $30.29$ $60.52$ $65.32$ $57.33$ $71.82$ $61.8$
w3a16g128 OWC $3.68$ $55.38$ $12.5$ $2.61$ $31.66$ $60.14$ $66.12$ $53.19$ $71.71$ $58.88$
GPTQ $4.02$ $50.04$ $14.64$ $3.15$ $32.42$ $61.91$ $66.64$ $56.05$ $73.01$ $59.75$
CD $3.38$ $52.01$ $9.62$ $2.21$ $31.23$ $63.13$ $65.9$ $55.98$ $72.47$ $59.91$
BCD(k=2) $3.1$ $50.95$ $9.36$ $1.92$ $31.14$ $62.25$ $65.41$ $56.16$ $72.91$ $58.96$
OWC-CD $3.6$ $51.69$ $9.33$ $2.31$ $31.83$ $61.74$ $65.02$ $56.23$ $72.2$ $59.43$
OWC-CD + CD $3.82$ $50.39$ $9.5$ $1.77$ $32.51$ $62.25$ $66.33$ $56.91$ $71.27$ $60.06$
w4a16 OWC $4.02$ $46.96$ $10.63$ $2.66$ $33.19$ $61.99$ $66.7$ $59.26$ $73.67$ $61.72$
GPTQ $4.46$ $48.69$ $18.84$ $3.1$ $34.47$ $62.96$ $64.77$ $60.09$ $73.61$ $61.09$
CD $4.13$ $48.88$ $16.64$ $3.15$ $33.53$ $63.89$ $66.79$ $60.31$ $73.78$ $61.4$
BCD(k=2) $4.24$ $48.31$ $15.73$ $3.3$ $33.28$ $63.34$ $66.39$ $60.26$ $73.56$ $61.01$
w4a16g128 OWC $4.99$ $46.1$ $20.62$ $4.28$ $34.73$ $64.48$ $67.83$ $59.91$ $74.32$ $61.25$
GPTQ $5.18$ $49.79$ $18.17$ $3.15$ $35.15$ $64.6$ $66.24$ $60.21$ $74.32$ $61.88$
CD $5.73$ $46.96$ $24.17$ $4.53$ $36.18$ $66.04$ $67.8$ $60.7$ $73.67$ $61.25$
BCD(k=2) $5.79$ $47.28$ $24.11$ $4.58$ $36.01$ $65.7$ $67.06$ $60.73$ $73.61$ $61.01$
OWC-CD $5.54$ $47.24$ $21.71$ $4.72$ $34.3$ $64.73$ $68.01$ $60.72$ $73.56$ $61.01$
OWC-CD + CD $5.62$ $47.6$ $24$ $4.68$ $35.49$ $66.16$ $67.06$ $60.78$ $74.1$ $61.48$
Table 16: Table presents downstream evaluation (zero-shot) numbers for GPTQ, CD, BCD for INT2 quantization of FFN weights. Gen, Rank columns correspond to generation and ranking tasks.
Method Epochs PaLM2-Otter PaLM2-Bison
w2a16g128 Generation Rank Avg. Generation Rank Avg.
GPTQ - $23.06$ $66.62$ $49.2$ $33.62$ $72.74$ $57.09$
CD 1 $25.12$ $65.73$ $49.48$ $33.63$ $71.80$ $56.53$
BCD(k=2) $1$ $25.42$ $65.57$ $49.51$ $33.97$ $72.16$ $56.88$
$2$ $24.73$ $65.78$ $49.36$ $34.01$ $72.39$ $57.04$
$3$ $25.28$ $66.01$ $49.72$ $34.27$ $72.67$ $57.31$
BCD(k=4) $1$ $25.15$ $66.04$ $49.68$ $33.94$ $72.10$ $56.84$
$2$ $24.86$ $65.71$ $49.37$ $33.87$ $71.97$ $56.73$
$3$ $25.70$ $65.30$ $49.46$ $33.85$ $72.06$ $56.78$
@@
Table 17: Table presents downstream evaluation (zero-shot) numbers of PaLM2-Otter - for GPTQ, CD, BCD for INT2 quantization of FFN weights.
Method Epochs PaLM2-Otter
@@
w2a16g128 NatualQ. SQuAD TriviaQA WebQ ARC-c ARC-e BoolQ HellaSwag PIQA WinoGrande
GPTQ - $4.32$ $63.43$ $22.08$ $2.41$ $42.32$ $66.96$ $80.09$ $68.54$ $73.78$ $68.03$
CD $1$ $4.99$ $66.51$ $25.53$ $3.44$ $40.87$ $70.33$ $74.89$ $68.13$ $72.96$ $67.17$
BCD(k=2) $1$ $4.93$ $67.51$ $25.56$ $3.69$ $41.47$ $69.7$ $74.98$ $67.97$ $73.72$ $65.59$
$2$ $4.6$ $67.74$ $23.6$ $3$ $40.78$ $69.49$ $75.47$ $68.27$ $73.39$ $67.25$
$3$ $5.1$ $67.91$ $24.53$ $3.59$ $42.41$ $69.87$ $75.02$ $68.52$ $73.88$ $66.38$
BCD(k=4) $1$ $4.9$ $67.5$ $25.01$ $3.2$ $42.15$ $68.69$ $76.18$ $68.51$ $73.61$ $67.09$
$2$ $4.79$ $67.9$ $23.55$ $3.2$ $41.81$ $69.82$ $73.39$ $68.5$ $73.5$ $67.25$
$3$ $5.46$ $68.29$ $26$ $3.05$ $41.21$ $69.02$ $73.98$ $68.06$ $72.52$ $67.01$
@@
Table 18: Table presents downstream evaluation (zero-shot) numbers of PaLM2-Bison - for GPTQ, CD, BCD for INT2 quantization of FFN weights.
Method Epochs PaLM2-Bison

### v3

@@ F.2 Detailed Downstream Evaluation Results
w16a16 - $12.85$ $65.27$ $58.34$ $8.46$ $51.79$ $76.09$ $81.8$ $79.3$ $79.6$ $72.22$
w3a16 OWC $2.27$ $49.45$ $8.02$ $0.49$ $41.98$ $60.14$ $48.78$ $66.91$ $71.87$ $66.06$
GPTQ $9.09$ $64.53$ $46.38$ $7.19$ $49.83$ $73.99$ $81.96$ $76.87$ $78.89$ $70.09$
CD $10.72$ $64.01$ $48.03$ $6.59$ $50.77$ $76.6$ $81.01$ $75.77$ $77.97$ $70.09$
BCD(k=2) $10.55$ $64.31$ $47.25$ $6.05$ $50.85$ $76.43$ $81.41$ $75.91$ $78.45$ $70.24$
GPTQ $10.17$ $66.53$ $49.29$ $7.73$ $51.37$ $75.93$ $82.35$ $77$ $79.11$ $72.14$
CD $7.17$ $68.73$ $49.42$ $5.61$ $50.26$ $76.18$ $82.39$ $76.79$ $78.13$ $71.98$
BCD(k=2) $6.68$ $68.18$ $48.39$ $5.27$ $50.26$ $75.93$ $81.41$ $76.84$ $79$ $72.61$
OWC-CD $8.98$ $68.37$ $48.33$ $5.91$ $50.51$ $75.46$ $81.35$ $76.83$ $78.18$ $72.06$
OWC-CD + CD $9.72$ $67.87$ $49.87$ $6.64$ $51.37$ $76.68$ $81.59$ $76.78$ $77.91$ $71.51$
w4a16 OWC $11.19$ $66.67$ $49.56$ $7.63$ $50$ $73.91$ $80$ $77.36$ $79.11$ $72.93$
GPTQ $12.27$ $66.65$ $48.72$ $8.46$ $51.71$ $75.84$ $82.02$ $78.2$ $80.09$ $73.32$
CD $12.52$ $66.85$ $53.16$ $7.53$ $52.3$ $76.43$ $81.28$ $78.3$ $79.6$ $72.61$
BCD(k=2) $12.33$ $67$ $53.45$ $7.63$ $51.96$ $75.97$ $80.92$ $78.47$ $80.47$ $72.06$
GPTQ $11.5$ $65.81$ $56.36$ $7.68$ $51.71$ $76.52$ $82.08$ $78.86$ $79.27$ $73.16$
CD $11.72$ $66.85$ $57.63$ $8.61$ $51.19$ $76.6$ $82.51$ $78.89$ $79.49$ $73.32$
BCD(k=2) $11.66$ $66.73$ $57.5$ $8.56$ $51.19$ $76.3$ $82.29$ $78.88$ $79.27$ $73.32$
OWC-CD $12.27$ $66.28$ $58.21$ $8.56$ $51.62$ $76.85$ $82.32$ $78.85$ $79.49$ $72.61$
OWC-CD + CD $11.63$ $66.28$ $58.06$ $7.92$ $51.62$ $76.94$ $82.45$ $78.72$ $79.11$ $72.69$
@@
Table 14: Table presents downstream evaluation (zero-shot) numbers of Gemma-1 7B - for GPTQ, CD, BCD for INT3, INT4 quantization of FFN weights.
Config Method Gemma-1 7B
@@
NaturalQ SQuAD TriviaQA WebQ ARC-c ARC-e BoolQ HellaSwag PIQA WinoGrande
w16a16 $15.04$ $71.94$ $62.38$ $13.04$ $43.69$ $65.53$ $80.43$ $78.55$ $80.79$ $72.14$
w316 OWC $0.03$ $2.87$ $0.15$ 00 $30.55$ $38.8$ $63.09$ $29$ $52.61$ $48.86$
GPTQ $2.24$ $58.68$ $14.35$ $3.05$ $35.15$ $53.16$ $63.76$ $53.84$ $64.58$ $59.91$
CD $6.04$ $68.77$ $32.22$ $4.28$ $41.04$ $62.37$ $76.51$ $68.95$ $74.76$ $67.01$
BCD(k=2) $6.76$ $69.03$ $33.32$ $3.74$ $40.7$ $62.21$ $76.91$ $69.1$ $75.24$ $68.11$
w3a16g128 OWC $6.51$ $71.04$ $30.86$ $5.46$ $38.14$ $60.1$ $70.7$ $64.51$ $75.3$ $63.85$
GPTQ $10.03$ $71.66$ $42.25$ $7.82$ $41.64$ $64.06$ $80.83$ $72.37$ $78.18$ $69.38$
CD $10.08$ $71.94$ $46.39$ $8.46$ $40.7$ $62.75$ $80.43$ $73.88$ $78.56$ $67.64$
BCD(k=2) $10.78$ $71.99$ $46.65$ $8.66$ $42.32$ $63.01$ $81.25$ $74.17$ $78.18$ $69.46$
OWC-CD $9.7$ $71.62$ $41.44$ $8.27$ $41.47$ $64.44$ $79.94$ $72.19$ $77.53$ $67.32$
OWC-CD + CD $10.66$ $71.78$ $46.73$ $8.02$ $42.41$ $64.73$ $80.46$ $74.29$ $78.94$ $67.64$
OWC-CD + BCD(k=2) $10.69$ $71.84$ $47.05$ $8.76$ $42.41$ $65.36$ $80.03$ $74.47$ $79.49$ $68.9$
w4a16 OWC $7.09$ $63.51$ $44.98$ $7.19$ $41.38$ $59.34$ $77.06$ $67.29$ $73.29$ $63.85$
GPTQ $10.66$ $70.95$ $52.63$ $9.55$ $44.11$ $64.56$ $79.39$ $75.9$ $79.16$ $70.56$
CD $12.71$ $70.41$ $57.39$ $11.27$ $43.69$ $65.11$ $79.66$ $77.02$ $79.98$ $70.17$
BCD(k=2) $12.44$ $70.5$ $57.55$ $11.96$ $45.14$ $65.4$ $79.79$ $76.66$ $79.54$ $71.03$
w4a16g128 OWC $13.55$ $71.61$ $56.09$ $12.8$ $42.49$ $63.85$ $80.92$ $76.77$ $79.22$ $70.32$
GPTQ $13.88$ $71.94$ $59.45$ $14.12$ $43.52$ $66.12$ $81.96$ $77.54$ $79.98$ $71.11$
CD $15.26$ $71.73$ $59.81$ $13.24$ $45.05$ $65.45$ $81.25$ $77.52$ $80.58$ $70.88$
BCD(k=2) $15.54$ $71.94$ $59.83$ $13.24$ $43.69$ $64.6$ $80.89$ $77.68$ $81.23$ $71.43$
OWC-CD $14.43$ $72.08$ $58.69$ $12.55$ $43.52$ $64.94$ $80.46$ $77.46$ $79.49$ $70.64$
OWC-CD + CD $14.82$ $72.32$ $59.98$ $13.63$ $45.05$ $65.82$ $80.43$ $77.46$ $80.96$ $71.43$
OWC-CD + BCD(k=2) $14.93$ $72.39$ $59.49$ $12.94$ $44.45$ $66.08$ $80.21$ $77.6$ $80.03$ $70.8$
@@
Table 15: Table presents downstream evaluation (zero-shot) numbers of Gemma-2 9B - for GPTQ, CD, BCD for INT3, INT4 quantization of FFN weights.
Config Method Gemma-2 9B